# Euclid Quick Data Release (Q1) point-like X-ray selected sources

### v1

@@ 3.1 XMM-Newton
The XMM-Newton space-telescope provides some of the deepest and most detailed X-ray imaging available, with a spatial resolution of approximately 6″ and positional uncertainties typically below 1 – 2″. The data used in this study include both targeted and serendipitous observations, selected from the 4XMM-DR14 catalogue333catalogue and data model description available at https://cdsarc.cds.unistra.fr/viz-bin/cat/IX/69, which offers multi-band flux measurements and quality flags. To prepare the sample, point-like sources were filtered based on their extent likelihood and positional uncertainties: SC_EXT_ML 10, SC_EXTENT 10 and SC_POSERR 10″, where the threshold for the positional uncertainty is motivated by the respective distribution as a function of flux as shown in Fig. 2. This effectively reduces the computational load for cross-matching with Euclid. To be consistent with other X-ray surveys, we compute the X-ray flux for the 0.5 – 2 keV band as
@@
by combining the 0.5 – 1 keV, and 1 – 2 keV band fluxes and errors.
@@ 3.2 Chandra
The Chandra X-ray space-telescope provides sub-arcsecond angular resolution, which allows precise localisation of sources and makes it particularly effective in crowded or complex fields. Our sample incorporates sources from the CSC 2.0 catalogue444https://vizier.cds.unistra.fr/viz-bin/VizieR?-source=IX/57&-to=3. To subsample only point-like sources, we select sources with the extended flag, , set to zero. The 0.5 – 2 keV X-ray flux and its uncertainty are computed as
with Flux and Flux referring to the default CSC 2.0 soft and medium bands, respectively, while (B) and (b) represent the upper and lower 1 error margins. The positional uncertainties are determined using the 2 errors along the semi-major (0) and semi-minor (1) axes of the error ellipse provided in the source catalogue, as
and downsample the CSC 2.0 catalogue such that C_POSERR 10″, to be consistent with 4XMM DR14 (see Fig. 2).
@@
The eROSITA-DE DR1 main catalogue555https://erosita.mpe.mpg.de/dr1/AllSkySurveyData_dr1/Catalogues_dr1/ offers comprehensive coverage of the western galactic hemisphere. While it does not include the EDF-N region (see Fig. 1), with a half energy width (HEW) of 26″ and a soft X-ray energy range of 0.2 – 2.3 keV, eROSITA offers a uniform data set ideal for large statistical studies. The data preparation for eROSITA included filtering for reliable point-source detections with EXT_LIKE 0. Additionally, we required that all nine quality flags (QFs) are set to zero, and applied a cut in positional uncertainty of POS_ERR 20″(refer to Footnote 2).
@@ 4 Identification of CTPs
Until recently, the identification of CTPs was primarily performed using maximum likelihood matching (e.g., Sutherland & Saunders 1992). This approach considers the separation between sources in the primary catalogue (e.g., X-ray) and the secondary multiwavelength catalogue, the positional uncertainties of the primary sources, and the magnitude distribution of sources in the ancillary catalogues within a certain radius from the primary sources. While this method has been applied successfully in cases with small positional errors and bright sources (see e.g., Naylor et al. 2013; Brusa et al. 2010; Xue et al. 2011b), its reliability decreases in situations where positional errors are large, and more than one photometric point is necessary for identifying the correct CTP or when multiple potential CTPs exist within the search radius. In such cases, positional matching often fails to provide a robust identification.
@@ 4.1 NWAY
To address these challenges, bayesian statistic needs to be invoked (Budavári & Szalay 2008). With this in mind, algorithms such as NWAY (Salvato et al. 2018), and XMATCH (Pineau et al. 2017) were developed. Unlike the others, NWAY also allows for the adoption of priors in addition to positional matches between multiple catalogues. It refines its matching procedure by being able to supplement one or more features, such as colours, magnitudes, and S/N, into likelihood ratios used for CTP identification. These priors, based on (anti-) correlations between source properties of the target and field populations, allow NWAY to improve the likelihood of finding true matches, boosting the accuracy significantly over methods that rely solely on positional information.
@@ 4.1.1 The integration of priors in NWAY
Colour and magnitude priors are posteriors that encode knowledge from data or probabilistic distributions that describe the expected properties of true CTPs for the sources in the primary catalogue. For example, an AGN selected in X-rays is likely to have distinct colours (e.g., redder MIR colours due to dust) and magnitudes compared to stars or inactive galaxies. The difference between the methods resides then in the adoption of specific features able to distinguish an X-ray emitter (regardless of its Galactic or extragalactic nature) from a random source in the field (Salvato et al. 2022). NWAY uses these priors to add additional dimensions to calculate a more nuanced likelihood ratio. This reduces ambiguity, especially in crowded fields or cases with large positional uncertainties.
For instance, in the ROSAT All-Sky Survey (2RXS, Boller et al. 2016), Salvato et al. (2018) improved CTP identification by incorporating a 2D prior based on the W2 magnitude and W1W2 colour from AllWISE (Wright et al. 2010), effectively separating out AGN. In the more recent application to the eROSITA Final Equatorial-Depth Survey (eFEDS, Brunner et al. 2022), Salvato et al. (2022) defined a prior that utilised not only magnitudes and colours from the Legacy Survey Data Release 10 (Dey et al. 2019), but also S/N in all bands (griz, W1, W2, W3, and W4), as well as properties, such as proper motion, from Gaia DR3 (Vallenari et al. 2023) when available. For further details on the formalism, we refer to the NWAY documentation777https://github.com/JohannesBuchner/NWAY.
@@ 4.1.2 Q1 prior
In this work, we have adopted the same principle as described in the previous subsection, but restricted ourselves to using only photometric features from Euclid. To build priors based on multiple features, NWAY can incorporate probabilities derived from ML models, such as a random forest (RF) classifier (the sklearn implementation, Pedregosa et al. 2011). We start by utilising a sample of secure CTPs to X-ray sources from 4XMM DR11 (Traulsen et al. 2020) and CSC 2.0 for training the RF as presented by Salvato et al. (2022). NWAY is then applied to identify CTPs for these X-ray sources in the EDF catalogues. A search radius of three times the positional error threshold is set around each X-ray source to include all potential EDF matches within this radius. This large search radius ensures that widely separated CTPs are not missed, even for sources with the largest positional uncertainties. The sky coverage of each X-ray sample per EDF is calculated to account for overlapping search windows. Specifically, while the coverage for the X-ray sample is given by its multi-order coverage map (MOC) or pointing, the EDF coverage is derived as
@@
where corresponds to the number of X-ray sources and is the overlapping area of adjacent search windows. Following the methodology outlined in Salvato et al. (2018), these coverage areas are used to compute the number densities, which inform two key probabilities: p_any (the probability that an X-ray source has a CTP) and p_i (the probability that each EDF source within the search radius is the correct CTP). NWAY identifies the most likely CTP for each X-ray source by selecting the EDF source with the highest p_i value. This CTP is considered the primary match. To assess the reliability of this identification, the p_any value is used as a confidence threshold. All sources identified as primary matches (match_flag == 1) with p_any ¿ 0.85, form the target sample. In addition, for every X-ray source, all other EDF sources within the search radius that are not the primary match are collected into a field population, provided p_any ¡ 0.1. This field population serves as a representation of non-X-ray sources, providing a comparison set for training the RF. To label the data sets, X-ray CTPs are assigned a target class of ’1’, while field objects are labelled as ‘0‘. The Euclid-only training features utilised for the RF classifier are outlined in Footnote 6. At present, photometry and colours from different apertures are not included as features, as the aperture sizes vary between sources despite sharing the same designation.
@@
Future data releases may standardise aperture sizes across all objects, also enhancing the RF classifier’s ability to leverage the relationship between light profiles and X-ray emission. To evaluate model performance, 15% of the training data set (6000 field and 1730 target sources) is randomly set aside as a test set. The RF model is configured with 200 decision trees, with splits allowed if at least 25 samples remain in a branch. Per decision tree construction, 10 photometric features are considered and bootstrap sampling is applied to the training set for building the ensemble of trees. Since the data set is highly imbalanced, with field objects vastly outnumbering CTPs of X-ray sources, a weighting scheme is implemented to automatically adjust the contribution of each class during training (Salvato et al. 2022). This ensures the model effectively learns to differentiate between the two classes despite the imbalance. The classifier outputs a probability score X-ray, the predicted likelihood of a candidate CTP being X-ray emitting. is used for class prediction, where NWAY retains candidates with if their astrometric configuration strongly supports a match. Importantly, the probability is derived exclusively from photometric properties, independent of positional uncertainties, enabling it to be seamlessly integrated into NWAY’s Bayesian framework, complementing astrometric priors with photometric information.
@@
We quantify the classification quality by comparing the input labels and predicted classes as: (i) true positive (TP, where label 1 and prediction 1); (ii) true negative (TN, where label 0 and prediction 0); (iii) false positive (FP, where label 0 and prediction 1); and (iv) false negative (FN, where label 1 and prediction 0). The model’s performance is summarised in the confusion matrix shown in Fig. 3, assessed by acquiring these values for the test set (see e.g., Fotopoulou & Paltani 2018). We define the following measures of quality for the classifier:
@@
The trained model achieves strong performance, with high accuracy of 89% and recall of 91%, indicating that most real X-ray emitters are correctly identified. The contamination rate remains low, with a fractional leakage of 11%. The fall-out fraction can, in principle, be reduced by including more field objects, though at the price of reduced recall.
@@ 4.2 Adding the X-ray emitter prior to NWAY
The a-priori probability of an Euclid source to be an X-ray emitter, , is now incorporated as a prior into the NWAY framework to enhance our CTP identification process. To visualise the positive impact of the inclusion of this prior, we plot the distribution of p_any both before and after the inclusion of the X-ray emitter prior (see Fig. 4). Incorporating the prior shifts the distribution of p_any to higher values, making the identification of true CTPs more reliable. Notably, the inclusion of the prior also changes the respective CTP in about 30% of the cases, indicating that for those sources that are indeed not the CTP after adding the prior, the p_any value has been decreased.
In total, we associate 12 645 CTP candidates. Among these, 11 286 (about ) sources have , indicating that they have the highest probability p_i to be considered reliable CTPs. The remaining candidates, with , represent secondary but plausible matches that are not ruled out entirely. We find 1092 X-ray sources with multiple CTPs, the vast majority of which is made up by XMM-Newton and eROSITA with only Chandra providing the required positional accuracy for Euclid to have few sources with multiple plausible CTPs. While most of the multi-CTP cases are made up of two candidates, a few cases with CTPs are introduced, predominantly by eROSITA, given its larger positional uncertainty. Notably, only five X-ray sources within the Q1 footprint (see Footnote 2) lack an association entirely, possibly due to boundary effects where the true CTP falls outside the region covered by the EDFs.
With this catalogue, several scenarios emerge in terms of the relationships between CTPs and X-ray sources.
@@
Two X-ray entries from differing surveys are assigned to a single CTP. This scenario suggests but does not confirm the possibility of a single physical source being detected by two different X-ray surveys.
@@
Cases where a single Euclid ID is linked to multiple entries from e.g., two or even all three X-ray surveys.
@@
Among the surveys, only eROSITA uniformly covers two of the three fields. In contrast, other surveys consist of pointed observations with differing depths and catalogues across distinct bands, meaning sources are not always detected at the same flux levels. As a result, the overlap between catalogues is complex and interpreting source matches becomes challenging. At this stage, gaps in coverage further limit the significance of overlap checks. Keeping this in mind we find 674 pairs which either correspond to case one or two, as well as 16 triplets referring to case three as shown in Fig. 5. We deliberately choose not to quantify case four, as there is currently no definitive way to resolve it with certainty. A breakdown of the CTP sample is given in Footnote 8.
@@ 4.3 Purity and completeness
To calibrate a reliable p_any cut-off, we perform the NWAY procedure (using the same prior) on a duplicate sample of the X-ray sources, referred to as ’randoms‘. These randoms are generated by displacing the positions of the X-ray sources to lie beyond the search radius, ensuring that any matches with multiwavelength CTPs are purely coincidental. By comparing the distribution of p_any values for both the original X-ray sample and the randoms, we can assess the impact of the prior by examining the two histograms in Fig. 4. The figure shows how the p_any value for the randoms is usually lower than for the real sample of X-ray sources, indicating that in a random position in the sky, there are no sources that have the same features of a typical X-ray emitters. For the few sources for which p_any remains high, either the prior was not representative enough, or, deeper data in the future will potentially reveal a fainter X-ray source.
@@
To further assess the impact of incorporating into the NWAY procedure, we compare the p_any values as a function of the separation distance between the primary (X-ray) and secondary (multi-wavelength) sources. Figure 6 presents the separation between the 4XMM sources and the associated CTP in EDF-F as a function of p_any. The sources with a higher probability of being correct also have a high probability of being X-ray emitters, even at larger separation, the latter being mostly smaller than 10″. CTPs of little separation and low p_any showcase scenarios of chance alignments where objects appear close but do not show significant . Next, we focus on evaluating the purity and completeness of the CTP samples based on how they are sub-selected. In this context, we define completeness as the cumulative fraction of sources with p_any higher than a certain value and purity the fraction of sources with p_any higher than the value in the random sample. This gives a measure of how often the CTP in the real sample is associated to a CTP only by chance.
@@
Figure 7 shows both purity and completeness for each EDF and X-ray survey as well as the ’sweet spot’ for the best trade off between these quantities as a function of p_any. Variations in the distributions result from differences in the positional uncertainties of the X-ray surveys and the source densities within both the X-ray data and the corresponding EDF. Depending on the scientific goal, a user interested in creating a catalogue that is extremely complete or pure, is hence free to select sources with a p_any beyond a certain threshold. For the rest of the work, we adopt the intersection between the purity and completeness curves as the threshold. In all fields, purity and completeness are at least 80%, with the Chandra catalogues unsurprisingly having the most reliable associations, reaching values of about 90%. Interestingly, the impact of the higher spatial resolution of Chandra is clear in the EDF-F field, where Fornax, the second richest galaxy cluster of the local Universe is located. In such a dense environment, it is more challenging to identify the correct CTP given the resolution of 4XMM and eROSITA.
@@
To conclude this part, we assess the stability in the method by examining the impact of RF hyper-parameters like minimum leaf size and features per tree. Variations in these parameters mainly shift the intersection of purity and completeness curves along the p_any axis, with smaller leaf sizes enhancing subclass specificity and subtly altering the intersection point.

### v2

@@ 3.1 XMM-Newton
The XMM-Newton space-telescope provides some of the deepest and most detailed X-ray imaging available, with a spatial resolution of approximately 6″ and positional uncertainties typically below 1–2″. The data used in this study includes both targeted and serendipitous observations, selected from the 4XMM DR14 catalogue333catalogue and data model description available at https://cdsarc.cds.unistra.fr/viz-bin/cat/IX/69, the most recent of the XMM releases444We note that dedicated XMM observations targeting the Fornax field are ongoing. These additional data are not included in this work but may be used to refine and extend the CTPs in future.. The catalogue offers multi-band flux measurements and quality flags. To prepare the sample, point-like sources were filtered based on their extent likelihood and positional uncertainties: SC_EXT_ML 10, SC_EXTENT 10 and SC_POSERR 10″, where the threshold for the positional uncertainty is motivated by the respective distribution as a function of flux as shown in Fig. 2. This effectively reduces the computational load for cross-matching with Euclid. To be consistent with other X-ray surveys, we computed the X-ray flux, , for the 0.5–2 keV band as
@@
by combining the 0.5–1 keV and 1–2 keV band fluxes and errors.
@@ 3.2 Chandra
The Chandra X-ray space telescope provides sub-arcsecond angular resolution, which allows for precise localisation of sources and makes it particularly effective in crowded or complex fields. Our sample incorporates sources from the CSC 2.0 catalogue555https://vizier.cds.unistra.fr/viz-bin/VizieR?-source=IX/57&-to=3. To subsample only point-like sources, we selected sources with the extended flag, , set to zero. The 0.5–2 keV X-ray flux and its uncertainty were computed as
with Flux and Flux referring to the default CSC 2.0 soft and medium bands, respectively, while (B) and (b) represent the upper and lower 1 error margins. The positional uncertainties were determined using the 2 errors along the semi-major (0) and semi-minor (1) axes of the error ellipse provided in the source catalogue as
and we downsampled the CSC 2.0 catalogue such that C_POSERR 10″to be consistent with 4XMM DR14 (see Fig. 2).
@@
The eROSITA-DE DR1 main catalogue666https://erosita.mpe.mpg.de/dr1/AllSkySurveyData_dr1/Catalogues_dr1/ offers comprehensive coverage of the western galactic hemisphere. While it does not include the EDF-N region (see Fig. 1), with a half energy width (HEW) of 26″ and a soft X-ray energy range of 0.2–2.3 keV, eROSITA offers a uniform data set ideal for large statistical studies. The data preparation for eROSITA included filtering for reliable point-source detections with EXT_LIKE 0. Additionally, we required that all nine quality flags (QFs) are set to zero, and applied a cut in positional uncertainty of POS_ERR 20″(refer to Footnote 2).
@@ 4 Identification of CTPs
Until recently, the identification of CTPs was primarily performed using maximum likelihood matching (e.g., Sutherland & Saunders 1992). This approach considers the separation between sources in the primary catalogue (e.g. X-ray) and the secondary multiwavelength catalogue, the positional uncertainties of the primary sources, and the magnitude distribution of sources in the ancillary catalogues within a certain radius from the primary sources. While this method has been applied successfully in cases with small positional errors and bright sources (see e.g., Naylor et al. 2013; Brusa et al. 2010; Xue et al. 2011b), its reliability decreases in situations where positional errors are large, and more than one photometric point is necessary for identifying the correct CTP or when multiple potential CTPs exist within the search radius. In such cases, positional matching often fails to provide a robust identification.
@@ 4.1 NWAY
To address these challenges, Bayesian statistics needs to be invoked (Budavári & Szalay 2008). With this in mind, algorithms such as NWAY (Salvato et al. 2018), and XMATCH (Pineau et al. 2017) were developed. Unlike the others, NWAY also allows for the adoption of priors in addition to positional matches between multiple catalogues. It refines its matching procedure by being able to supplement one or more features, such as colours, magnitudes, and S/N, into likelihood ratios used for CTP identification. These priors based on (anti-) correlations between source properties of the target and field populations allow NWAY to improve the likelihood of finding true matches, boosting the accuracy significantly over methods that rely solely on positional information.
@@ 4.1.1 The integration of priors in NWAY
Colour and magnitude priors are posteriors that encode knowledge from data or probabilistic distributions that describe the expected properties of true CTPs for the sources in the primary catalogue. For example, an AGN selected in X-rays is likely to have distinct colours (e.g. redder MIR colours due to dust) and magnitudes compared to stars or inactive galaxies. The difference between the methods resides then in the adoption of specific features able to distinguish an X-ray emitter (regardless of its Galactic or extragalactic nature) from a random source in the field (Salvato et al. 2022). NWAY uses these priors to add additional dimensions to calculate a more nuanced likelihood ratio. This reduces ambiguity, especially in crowded fields or cases with large positional uncertainties.
For instance, in the ROSAT All-Sky Survey (2RXS, Boller et al. 2016), Salvato et al. (2018) improved CTP identification by incorporating a 2D prior based on the W2 magnitude and W1W2 colour from AllWISE (Wright et al. 2010), effectively separating out AGN. In the more recent application to the eROSITA Final Equatorial-Depth Survey (eFEDS, Brunner et al. 2022), Salvato et al. (2022) defined a prior that utilised not only magnitudes and colours from the Legacy Survey Data Release 10 (LS10, Dey et al. 2019), but also S/N in all bands (griz, W1, W2, W3, and W4), as well as properties, such as proper motion, from Gaia DR3 (Vallenari et al. 2023) when available. For further details on the formalism, we refer to the NWAY documentation.888https://github.com/JohannesBuchner/NWAY
@@ 4.1.2 Q1 prior
In this work, we have adopted the same principle as described in the previous subsection, but restricted ourselves to using only photometric features from Euclid. To build priors based on multiple features, NWAY can incorporate probabilities derived from ML models, such as a random forest (RF) classifier (the sklearn implementation, Pedregosa et al. 2011). We start by utilising the training sample presented in Salvato et al., in prep., comprising secure CTPs to X-ray sources from 4XMM DR11 (Webb et al. 2020) and CSC 2.0, for training the RF model, as presented by (Salvato et al. 2022). NWAY is then applied to identify CTPs for these X-ray sources in the EDF catalogues. A search radius of three times the positional error threshold is set around each X-ray source to include all potential EDF matches within this radius. This large search radius ensures that widely separated CTPs are not missed, even for sources with the largest positional uncertainties. The sky coverage of each X-ray sample per EDF is calculated to account for overlapping search windows. Specifically, while the coverage for the X-ray sample is given by its multi-order coverage map (MOC) or pointing, the EDF coverage was derived as
@@
where corresponds to the number of X-ray sources and is the overlapping area of adjacent search windows. Following the methodology outlined in Salvato et al. (2018), these coverage areas are used to compute the number densities, which inform two key probabilities: p_any (the probability that an X-ray source has a CTP) and p_i (the probability that each EDF source within the search radius is the correct CTP). NWAY identifies the most likely CTP for each X-ray source by selecting the EDF source with the highest p_i value. This CTP is considered the primary match. To assess the reliability of this identification, the p_any value is used as a confidence threshold. All sources identified as primary matches (match_flag == 1) with p_any ¿ 0.85, form the target sample. In addition, for every X-ray source, all other EDF sources within the search radius that are not the primary match are collected into a field population, provided p_any ¡ 0.1. This field population serves as a representation of non-X-ray sources, providing a comparison set for training the RF. To label the data sets, X-ray CTPs are assigned a target class of ‘1’, while field objects are labelled as ‘0‘. The Euclid-only training features utilised for the RF classifier are outlined in Footnote 7. At present, photometry and colours from different apertures are not included as features, as the aperture sizes vary between sources despite sharing the same designation.
@@
Future data releases may standardise aperture sizes across all objects, also enhancing the RF classifier’s ability to leverage the relationship between light profiles and X-ray emission. To evaluate model performance, 15% of the training dataset (6000 fields and 1730 target sources) is randomly set aside as a test set. The RF model is configured with 200 decision trees, with splits allowed if at least 25 samples remain in a branch. Per decision tree construction, 10 photometric features are considered, and bootstrap sampling is applied to the training set for building the ensemble of trees. Since the data set is highly imbalanced, with field objects vastly outnumbering CTPs of X-ray sources, a weighting scheme is implemented to automatically adjust the contribution of each class during training (Salvato et al. 2022). This ensures the model effectively learns to differentiate between the two classes despite the imbalance. The classifier outputs a probability score X-ray, the predicted likelihood of a candidate CTP being X-ray emitting. is used for class prediction, where NWAY retains candidates with if their astrometric configuration strongly supports a match. Importantly, the probability is derived exclusively from photometric properties, independent of positional uncertainties, enabling it to be seamlessly integrated into NWAY’s Bayesian framework, complementing astrometric priors with photometric information.
@@
We quantify the classification quality by comparing the input labels and predicted classes as: (i) true positive (TP, where label 1 and prediction 1); (ii) true negative (TN, where label 0 and prediction 0); (iii) false positive (FP, where label 0 and prediction 1); and (iv) false negative (FN, where label 1 and prediction 0). The model’s performance is summarised in the confusion matrix shown in Fig. 3, assessed by acquiring these values for the test set (see e.g., Fotopoulou & Paltani 2018). We defined the following measures of quality for the classifier:
@@
The trained model achieves strong performance, with a high accuracy of 89% and a recall of 91%, indicating that most real X-ray emitters are correctly identified. The contamination rate remains low, with a fractional leakage of 11%. The fall-out fraction can, in principle, be reduced by including more field objects, though at the price of reduced recall.
@@ 4.2 Adding the X-ray emitter prior to NWAY
The a priori probability of a Euclid source being an X-ray emitter, , is now incorporated as a prior into the NWAY framework to enhance our CTP identification process. To visualise the positive impact of the inclusion of this prior, we plot the distribution of p_any both before and after the inclusion of the X-ray emitter prior (see Fig. 4). Incorporating the prior shifts the distribution of p_any to higher values, making the identification of true CTPs more reliable. Notably, the inclusion of the prior also changes the respective CTP in about 30% of the cases, indicating that for those sources that are indeed not the CTP after adding the prior, the p_any value has been decreased.
In total, we associated 12 645 CTP candidates. Among these, 11 286 sources (about ) have , indicating that they have the highest probability p_i to be considered reliable CTPs. The remaining candidates, with , represent secondary but plausible matches that are not ruled out entirely. We found 1092 X-ray sources with multiple CTPs, the vast majority of which are made up by XMM-Newton and eROSITA, with only Chandra providing the required positional accuracy for Euclid to have a few sources with multiple plausible CTPs. While most of the multi-CTP cases are made up of two candidates, a few cases with more than two CTPs are introduced, predominantly by eROSITA, given its larger positional uncertainty. Notably, only five X-ray sources within the Q1 footprint (see Footnote 2) lack an association entirely, possibly due to boundary effects where the true CTP falls outside the region covered by the EDFs.
With this catalogue, several scenarios emerged in terms of the relationships between CTPs and X-ray sources:
@@
Two X-ray entries from different surveys are assigned to a single CTP. This scenario suggests but does not confirm the possibility of a single physical source being detected by two different X-ray surveys.
@@
Cases where a single Euclid ID is linked to multiple entries from two or even all three X-ray surveys.
@@
Among the surveys, only eROSITA uniformly covers two of the three fields. In contrast, other surveys consist of pointed observations with differing depths and catalogues across distinct bands, meaning sources are not always detected at the same flux levels. As a result, the overlap between catalogues is complex and interpreting source matches becomes challenging. At this stage, gaps in coverage further limit the significance of overlap checks. Keeping this in mind, we find 674 pairs which either correspond to case one or two, as well as 16 triplets referring to case three, as shown in Fig. 5. We deliberately choose not to quantify case four, as there is currently no definitive way to resolve it with certainty. A breakdown of the CTP sample is given in Footnote 9.
@@ 4.3 Purity and completeness
To calibrate a reliable p_any cut-off, we performed the NWAY procedure (using the same prior) on a duplicate sample of the X-ray sources, referred to as ‘randoms’. These randoms were generated by displacing the positions of the X-ray sources to lie beyond the search radius, ensuring that any matches with multiwavelength CTPs are purely coincidental. By comparing the distribution of p_any values for both the original X-ray sample and the randoms, we could assess the impact of the prior by examining the two histograms in Fig. 4. The figure shows how the p_any value for the randoms is usually lower than for the real sample of X-ray sources, indicating that in a random position in the sky, there are no sources that have the same features as a typical X-ray emitter. For the few sources for which p_any remains high, either the prior was not representative enough, or, deeper data in the future will potentially reveal a fainter X-ray source.
@@
To further assess the impact of incorporating into the NWAY procedure, we compared the p_any values as a function of the separation distance between the primary (X-ray) and secondary (multi-wavelength) sources. Figure 6 presents the separation between the 4XMM sources and the associated CTP in EDF-F as a function of p_any. The sources with a higher probability of being correct also have a high probability of being X-ray emitters, even at larger separations, the latter being mostly smaller than 10″. CTPs of little separation and low p_any showcase scenarios of chance alignments where objects appear close but do not show significant . Next, we focused on evaluating the purity and completeness of the CTP samples based on how they are sub-selected. In this context, we defined completeness as the cumulative fraction of sources with p_any higher than a certain value and purity as the fraction of sources with p_any higher than the value in the random sample. This gives a measure of how often the CTP in the real sample is associated to a CTP only by chance.
@@
Figure 7 shows both the purity and completeness for each EDF and X-ray survey as well as the ‘sweet spot’ for the best trade-off between these quantities as a function of p_any. Variations in the distributions result from differences in the positional uncertainties of the X-ray surveys and the source densities within both the X-ray data and the corresponding EDF. Depending on the scientific goal, a user interested in creating a catalogue that is extremely complete or pure is hence free to select sources with a p_any beyond a certain threshold. For the rest of the work, we adopt the intersection between the purity and completeness curves as the threshold. In all fields, purity and completeness are at least 80%, with the Chandra catalogues unsurprisingly having the most reliable associations, reaching values of about 90%. Interestingly, the impact of the higher spatial resolution of Chandra is clear in the EDF-F field, where Fornax, the second richest galaxy cluster of the local Universe, is located. In such a dense environment, it is more challenging to identify the correct CTP given the resolution of 4XMM and eROSITA.
@@
To conclude this part, we assessed the stability of the method by examining the impact of RF hyperparameters such as the minimum leaf size and features per tree. Variations in these parameters mainly shift the intersection of purity and completeness curves along the p_any axis, with smaller leaf sizes enhancing subclass specificity and subtly altering the intersection point.